OpenStudy (anonymous):

what is integration of sinx-1/2(sin2x) frm 0 to pi

OpenStudy (turingtest):

\[\int_{0}^{\pi}\sin x-\frac12\sin(2x)dx\]?

OpenStudy (anonymous):

yes boss ,

OpenStudy (turingtest):

you know the integral of sinx I'm guessing, right?

OpenStudy (anonymous):

i need whole derivation

OpenStudy (turingtest):

the derivation that \[\int_a^b\sin xdx=-\cos x|_a^b\]?? that takes some doing...

OpenStudy (anonymous):

yp yp ,try if u can solve remaining part

OpenStudy (turingtest):

I'm quite sure I can, but you asked me to prove it the second integral can be done with the u-substitution\[u=2x\implies du=2dx\]

OpenStudy (anonymous):

thnx brother ,i shall try itknow ,thnx for help. ahmed